You are given an endpoint and the midpoint of a line segment. Find the coordinates of the other endpoint. Each midpoint is denoted by M(x, y)
Q (-1,14)

M (2,7)

Answer by HyperBrain(694)

Q (-1,14)
M (2,7)
---
Let A(h,k) be the unknown point.
---
Using the midpoint formula,
(h+(-1))/2=2
h+(-1)=4
h-1=4
h=5
---
Applying the midpoint formula for k,
(k+14)/2=7
k+14=14
k=0
---
So, the other endpoint is at (5,0)